# From Informal to Formal – Incorporating and Evaluating LLMs on Natural Language Requirements to Verifiable Formal Proofs

### v3

@@ 5 TLA+ “***Parse Error***”, “Unknown operator”
LLMs Size ProofGen SegGen ProofComplete ProofInfill Code2Proof
@@
llama3.1-instruct 8B 0/13 1/45 33/58 0/10 1/12 43/69 746/2910 78/124 4/33 4/63 4/11 45/81 16/34 1/5 16/34 44/157 27/60 0/13 1/7 0/4
llama3.1-instruct 70B 2/12 0/44 38/53 1/8 1/13 43/54 71/2802 75/114 16/32 5/38 6/8 44/81 10/31 1/5 15/34 39/153 22/49 0/13 2/11 3/11
qwen2.5-instruct 7B 10/13 1/46 30/56 1/10 0/11 26/69 303/2914 53/122 5/34 6/62 9/11 35/73 8/29 1/5 14/34 66/156 31/53 0/13 4/7 3/6
qwen2.5-instruct 72B 7/12 0/44 25/48 1/8 9/11 47/61 40/2848 54/116 11/27 0/55 7/11 35/74 8/24 1/5 13/34 59/149 20/49 0/11 0/9 1/6
qwen2.5-coder-instruct 7B 9/13 0/46 32/55 0/10 0/14 60/68 22/2882 72/116 6/32 53/61 10/11 12/80 8/32 2/6 17/34 34/159 26/50 1/12 0/7 0/6
qwen2.5-coder-instruct 32B 7/13 0/44 26/52 1/10 2/11 50/61 31/2825 68/119 9/27 25/46 5/8 24/79 6/23 3/4 17/33 51/152 16/43 0/12 0/4 0/6
deepseek-coder-instruct 7B 10/13 0/46 29/55 0/10 0/13 35/67 9/2894 62/127 3/34 43/63 9/10 14/82 11/31 2/5 22/34 18/160 23/56 0/10 1/13 1/12
deepseek-coder-instruct 33B 12/13 0/47 43/58 4/10 0/14 66/68 52/2888 81/126 8/32 45/61 10/11 3/81 6/31 2/4 19/34 6/161 23/50 0/14 0/11 1/13
starcoder-instruct 15B 8/12 0/46 24/54 2/7 3/10 36/51 10/2887 67/124 12/26 1/63 6/8 4/74 7/27 1/4 20/34 6/149 23/49 0/12 0/9 1/10
@@ Appendix E Complete Evaluation Result
ProofGen SegGen ProofComplete ProofInfill Code2Sepc
LLMs Size TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL
@@
llama3.1-instruct 8B 0.00 4.26 1.69 0.00 14.29 1.43 1.19 6.06 8.33 0.00 0.00 3.57 5.56 28.57 0.00 4.85 7.69 13.33 50.00 71.43
llama3.1-instruct 70B 7.69 6.38 10.17 20.00 7.14 22.86 4.86 13.64 11.11 39.68 27.27 3.57 13.89 28.57 0.00 7.27 24.62 13.33 21.43 21.43
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 21.43 1.43 1.05 7.58 5.56 1.59 0.00 13.10 19.44 28.57 0.00 5.45 18.46 13.33 50.00 57.14
qwen2.5-instruct 72B 7.69 6.38 18.64 20.00 21.43 12.86 3.29 12.12 25.00 12.70 0.00 11.90 33.33 28.57 0.00 9.70 24.62 26.67 35.71 57.14
qwen2.5-coder-instruct 7B 0.00 2.13 6.78 0.00 0.00 2.86 2.14 12.12 11.11 3.17 0.00 4.76 11.11 14.29 0.00 3.64 23.08 20.00 50.00 57.14
qwen2.5-coder-instruct 32B 0.00 6.38 11.86 0.00 21.43 12.86 4.07 9.85 25.00 26.98 27.27 5.95 36.11 42.86 2.94 7.88 33.85 20.00 71.43 57.14
deepseek-coder-instruct 7B 0.00 2.13 6.78 0.00 7.14 4.29 1.73 3.79 5.56 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 7.14 14.29
deepseek-coder-instruct 33B 0.00 0.00 1.69 0.00 0.00 2.86 1.94 4.55 11.11 3.17 0.00 3.57 13.89 42.86 0.00 2.42 23.08 6.67 21.43 7.14
star-coder-instruct 15B 7.69 2.13 8.47 30.00 28.57 27.14 1.97 6.06 27.78 0.00 27.27 11.90 25.00 42.86 0.00 9.70 24.62 20.00 35.71 28.57
@@
llama3.1-fma 8B 0.00 6.38 8.47 20.00 57.14 41.43 29.98 21.97 25.00 90.48 36.36 8.33 11.11 28.57 0.00 0.00 7.69 0.00 21.43 21.43
llama3.1-ultrachat 8B 0.00 0.00 1.69 0.00 7.14 0.00 0.00 3.79 0.00 0.00 0.00 5.95 13.89 14.29 0.00 4.85 9.23 0.00 14.29 0.00
llama3.1-ultrachat-fma 8B 0.00 8.51 10.17 30.00 50.00 41.43 35.72 29.55 33.33 95.24 18.18 8.33 11.11 28.57 2.94 0.00 7.69 0.00 28.57 35.71
llama3.1-tulu 8B 0.00 0.00 1.69 0.00 7.14 0.00 0.71 6.06 0.00 0.00 0.00 5.95 8.33 28.57 0.00 1.82 10.77 0.00 57.14 50.00
llama3.1-tulu-fma 8B 0.00 4.26 11.86 30.00 35.71 42.86 36.71 27.27 36.11 98.41 18.18 10.71 11.11 42.86 2.94 2.42 6.15 0.00 35.71 50.00
qwen2.5-fma 7B 0.00 4.26 11.86 10.00 71.43 38.57 27.95 27.27 22.22 87.30 27.27 9.52 13.89 28.57 0.00 0.00 7.69 0.00 35.71 21.43
qwen2.5-coder-fma 7B 0.00 6.38 18.64 20.00 35.71 44.29 36.50 34.09 33.33 98.41 36.36 9.52 16.67 28.57 5.88 1.82 13.85 13.33 42.86 50.00
deepseek-coder-fma 7B 0.00 2.13 16.95 0.00 21.43 34.29 25.30 31.06 25.00 84.13 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 21.43 35.71
@@
ProofGen SegGen ProofComplete ProofInfill Code2Sepc
LLMs Size TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL
@@
llama3.1-instruct 8B 0.00 2.13 5.08 0.00 35.71 4.29 1.90 8.33 11.11 3.17 0.00 3.57 8.33 28.57 0.00 7.88 20.00 20.00 71.43 85.71
llama3.1-instruct 70B 7.69 8.51 13.56 20.00 35.71 27.14 6.01 17.42 13.89 52.38 36.36 11.90 27.78 28.57 0.00 10.30 30.77 20.00 35.71 57.14
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 28.57 2.86 1.39 9.09 5.56 3.17 0.00 16.67 22.22 28.57 0.00 5.45 20.00 20.00 71.43 78.57
qwen2.5-instruct 72B 15.38 6.38 22.03 30.00 21.43 17.14 3.80 14.39 25.00 15.87 9.09 16.67 38.89 28.57 0.00 11.52 26.15 33.33 57.14 92.86
qwen2.5-coder-instruct 7B 0.00 4.26 10.17 0.00 7.14 5.71 2.99 13.64 11.11 4.76 0.00 11.90 19.44 28.57 0.00 4.24 30.77 20.00 85.71 92.86
qwen2.5-coder-instruct 32B 0.00 6.38 15.25 10.00 28.57 15.71 4.96 15.15 27.78 33.33 36.36 10.71 47.22 42.86 2.94 10.91 33.85 20.00 92.86 71.43
deepseek-coder-instruct 7B 7.69 2.13 8.47 0.00 7.14 8.57 2.82 5.30 5.56 4.76 18.18 9.52 16.67 28.57 0.00 4.24 18.46 33.33 57.14 64.29
deepseek-coder-instruct 33B 0.00 6.38 6.78 20.00 7.14 2.86 3.23 6.82 13.89 25.40 0.00 11.90 19.44 42.86 0.00 10.30 40.00 26.67 28.57 14.29
star-coder-instruct 15B 15.38 4.26 10.17 30.00 50.00 34.29 3.46 7.58 27.78 0.00 45.45 16.67 41.67 57.14 0.00 16.97 38.46 26.67 42.86 42.86
@@
llama3.1-fma 8B 0.00 6.38 15.25 20.00 57.14 41.43 34.57 26.52 27.78 90.48 36.36 9.52 11.11 28.57 2.94 0.00 13.85 0.00 42.86 57.14
llama3.1-ultrachat 8B 0.00 0.00 3.39 0.00 14.29 0.00 0.07 7.58 0.00 1.59 0.00 4.76 22.22 14.29 0.00 4.85 9.23 6.67 14.29 0.00
llama3.1-ultrachat-fma 8B 0.00 10.64 15.25 30.00 57.14 42.86 39.86 31.06 38.89 96.83 18.18 9.52 13.89 28.57 11.76 0.00 10.77 0.00 42.86 64.29
llama3.1-tulu 8B 0.00 2.13 1.69 0.00 21.43 0.00 1.43 6.06 0.00 3.17 0.00 13.10 30.56 28.57 0.00 6.67 21.54 13.33 78.57 64.29
llama3.1-tulu-fma 8B 0.00 6.38 11.86 30.00 35.71 44.29 40.20 32.58 38.89 98.41 18.18 13.10 11.11 42.86 2.94 4.24 9.23 0.00 71.43 64.29
qwen2.5-fma 7B 0.00 4.26 13.56 10.00 71.43 40.00 32.77 30.30 27.78 88.89 45.45 10.71 19.44 28.57 2.94 1.21 12.31 0.00 57.14 50.00
qwen2.5-coder-fma 7B 0.00 8.51 22.03 30.00 50.00 44.29 40.81 35.61 36.11 98.41 54.55 10.71 19.44 42.86 8.82 4.24 20.00 13.33 64.29 64.29
deepseek-coder-fma 7B 0.00 4.26 16.95 0.00 35.71 35.71 30.70 36.36 25.00 85.71 45.45 13.10 16.67 28.57 2.94 4.24 9.23 6.67 21.43 57.14
@@ Appendix F Prompt Design
Prompt for ProofComplete Task (lang: a placeholder to be replaced by each formal specification language name.) # Task Description (ProofInfill) 1. Please fill in the [MASK] in the following formal proof in formal specification language lang according to the given requirement. 2. Please fill in the [MASK] in the following formal proof in lang according to the given requirement. 3. Please complete the given formal proof in lang following the requirement below by filling in the [MASK]. 4. Please fill in the [MASK] in the following formal proof in lang according to the requirements below. 5. Please fill in the [MASK] in the following formal proof in lang according to the given requirement. (Randomly choose one of the above.) You only need to return the completed lang formal specification (together with the provided formal specification) without explanation. # Input
@@
Prompt for ProofComplete Task # Task Description (Code2Proof) 1. Please fill in the [MASK] in ACSL according to the given requirement and ACSL specification. 2. Please fill in the [MASK] in ACSL according to the given requirement. 3. Please fill in the [MASK] in ACSL according to the given ACSL specification. 4. Please fill in the [MASK] in ACSL according to the given requirement and ACSL specification. 5. Please infill the [MASK] in ACSL according to the given requirement. (Randomly choose one of the above.) You only need to return the completed ACSL formal specification (together with the provided formal specifications and C programs) without explanation. # Input

### v4

@@ 5 Conclusion
6 Acknowledgments
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported in part by the Hong Kong Research Grants Council/General Research Fund (HKSAR RGC/GRF, Grant No. 16206524), the National Natural Science Foundation of China (Grant Nos. 62372193, 62192734, 62302375, 62472339), CAS Project for Young Scientists in Basic Research (Grant No.YSBR-040) and the Basic Research Program of ISCAS (Grant No. ISCAS-ZD-202402).
@@ 5 TLA+ “***Parse Error***”, “Unknown operator”
LLMs Size ProofGen SegGen ProofComplete ProofInfill Cd2Prf
@@
llama3.1-instruct 8B 0/13 18/45 34/58 0/10 1/14 43/69 742/2910 78/124 4/33 4/63 4/11 45/81 16/34 1/5 16/34 44/157 27/60 0/13 1/14 0/14
llama3.1-instruct 70B 4/12 10/44 40/53 1/8 1/14 43/53 71/2802 75/114 16/32 5/61 6/8 44/81 10/31 1/5 15/34 39/153 22/49 0/13 2/14 3/14
qwen2.5-instruct 7B 12/13 13/46 33/56 1/10 0/14 26/69 302/2914 53/122 5/34 6/62 9/11 35/73 8/29 1/5 14/34 66/156 31/53 0/13 4/14 3/14
qwen2.5-instruct 72B 10/12 6/44 28/48 1/8 9/13 47/61 40/2848 54/116 11/27 0/60 7/11 35/74 8/24 1/5 13/34 59/149 20/49 0/11 0/14 1/14
qwen2.5-coder-instruct 7B 11/13 18/46 38/55 0/10 0/14 60/68 22/2882 72/116 6/32 53/63 10/11 12/80 8/32 2/6 17/34 34/159 26/50 1/12 0/14 0/14
qwen2.5-coder-instruct 32B 11/13 12/44 29/52 1/10 2/14 50/61 31/2825 68/119 9/27 25/60 5/8 24/79 6/23 3/4 17/33 51/152 16/43 0/12 0/12 0/12
deepseek-coder-instruct 7B 10/13 4/46 31/54 1/10 0/14 35/67 9/2894 62/127 3/34 43/63 9/10 14/82 11/31 2/5 22/34 18/160 23/56 0/10 1/14 1/14
deepseek-coder-instruct 33B 12/13 0/47 46/58 4/10 0/14 66/68 52/2888 81/126 8/32 45/63 10/11 3/81 6/31 2/4 19/34 6/161 23/50 0/14 0/14 1/14
star-coder-instruct 15B 8/12 0/46 27/54 2/7 3/14 36/51 10/2887 67/124 12/26 1/63 6/8 4/74 7/27 1/4 20/34 6/149 23/49 0/12 0/14 1/14
@@ Appendix E Complete Evaluation Result
ProofGen SegGen ProofComplete ProofInfill Cd2Prf
LLMs Size TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL-WP ACSL ACSL-WP
@@
llama3.1-instruct 8B 0.00 4.26 1.69 0.00 14.29 0.00 1.43 1.19 6.06 8.33 0.00 0.00 0.00 3.57 5.56 28.57 0.00 4.85 7.69 13.33 21.43 0.00 57.14 0.00
llama3.1-instruct 70B 7.69 6.38 10.17 20.00 7.14 0.00 24.29 4.86 13.64 11.11 39.68 3.17 27.27 3.57 13.89 28.57 0.00 7.27 24.62 13.33 21.43 0.00 21.43 0.00
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 21.43 0.00 1.43 1.05 7.58 5.56 1.59 1.59 0.00 13.10 19.44 28.57 0.00 5.45 18.46 13.33 21.43 0.00 14.29 0.00
qwen2.5-instruct 72B 7.69 6.38 18.64 20.00 21.43 7.14 12.86 3.29 12.12 25.00 12.70 4.76 0.00 11.90 33.33 28.57 0.00 9.70 24.62 26.67 35.71 0.00 57.14 0.00
qwen2.5-coder-instruct 7B 0.00 2.13 6.78 0.00 0.00 0.00 2.86 2.14 12.12 11.11 3.17 0.00 0.00 4.76 11.11 14.29 0.00 3.64 23.08 20.00 0.00 0.00 0.00 0.00
qwen2.5-coder-instruct 32B 0.00 6.38 11.86 0.00 21.43 0.00 12.86 4.07 9.85 25.00 26.98 4.76 27.27 5.95 36.11 42.86 2.94 7.88 33.85 20.00 71.43 14.29 57.14 14.29
star-coder-instruct 15B 7.69 2.13 8.47 30.00 28.57 0.00 27.14 1.97 6.06 27.78 0.00 0.00 27.27 11.90 25.00 42.86 0.00 9.70 24.62 20.00 35.71 0.00 28.57 0.00
deepseek-coder-instruct 7B 0.00 2.13 8.47 0.00 0.00 0.00 4.29 1.73 3.79 5.56 0.00 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 0.00 0.00 0.00 0.00
deepseek-coder-instruct 33B 0.00 0.00 1.69 0.00 0.00 0.00 2.86 1.94 4.55 11.11 3.17 0.00 0.00 3.57 13.89 42.86 0.00 2.42 23.08 6.67 7.14 0.00 7.14 0.00
deepseek-r1 xxB 30.77 8.51 30.51 30.00 35.71 7.14 22.86 10.70 21.21 22.22 49.21 3.17 45.45 15.48 33.33 28.57 5.88 9.70 24.62 6.67 42.86 21.43 64.29 14.29
@@
llama3.1-sft 8B 0.00 6.38 8.47 20.00 0.00 0.00 38.57 29.58 21.97 25.00 88.89 17.46 36.36 8.33 11.11 28.57 0.00 0.00 7.69 0.00 21.43 0.00 21.43 0.00
llama3.1-ultrachat 8B 0.00 6.38 3.39 0.00 0.00 0.00 0.00 0.00 3.79 0.00 0.00 0.00 0.00 5.95 13.89 14.29 0.00 4.85 9.23 0.00 14.29 7.14 0.00 0.00
llama3.1-ultrachat-fm 8B 0.00 8.51 10.17 30.00 0.00 0.00 41.43 35.79 29.55 33.33 95.24 17.46 18.18 8.33 11.11 28.57 2.94 0.00 7.69 0.00 28.57 0.00 35.71 0.00
llama3.1-tulu 8B 0.00 2.13 1.69 0.00 0.00 0.00 0.00 0.71 6.06 0.00 0.00 0.00 0.00 5.95 8.33 28.57 0.00 1.82 10.77 0.00 7.14 0.00 14.29 0.00
llama3.1-tulu-fm 8B 0.00 4.26 11.86 30.00 7.14 0.00 42.86 36.64 27.27 36.11 98.41 17.46 18.18 10.71 11.11 42.86 2.94 2.42 6.15 0.00 35.71 0.00 50.00 0.00
qwen2.5-sft 7B 0.00 4.26 11.86 10.00 0.00 0.00 38.57 27.95 27.27 22.22 87.30 17.46 27.27 9.52 13.89 28.57 0.00 0.00 7.69 0.00 21.43 0.00 14.29 0.00
qwen2.5-coder-sft 7B 0.00 6.38 18.64 20.00 7.14 0.00 44.29 36.50 34.09 33.33 98.41 17.46 36.36 9.52 16.67 28.57 5.88 1.82 13.85 13.33 42.86 0.00 42.86 0.00
deepseek-coder-sft 7B 0.00 2.13 16.95 0.00 0.00 0.00 34.29 25.30 31.06 25.00 84.13 15.87 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 14.29 7.14 14.29 7.14
@@
ProofGen SegGen ProofComplete ProofInfill Cd2Prf
LLMs Size TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP ACSL ACSL-WP
@@
llama3.1-instruct 8B 0.00 2.13 5.08 0.00 35.71 0.00 4.29 1.90 8.33 11.11 3.17 0.00 0.00 3.57 8.33 28.57 0.00 7.88 20.00 20.00 42.86 0.00 78.57 7.14
llama3.1-instruct 70B 7.69 8.51 13.56 20.00 35.71 21.43 27.14 6.01 17.42 13.89 52.38 4.76 36.36 11.90 27.78 28.57 0.00 10.30 30.77 20.00 35.71 7.14 50.00 14.29
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 28.57 0.00 2.86 1.39 9.09 5.56 3.17 3.17 0.00 16.67 22.22 28.57 0.00 5.45 20.00 20.00 42.86 0.00 42.86 0.00
qwen2.5-instruct 72B 15.38 6.38 22.03 30.00 21.43 7.14 17.14 3.80 14.39 25.00 15.87 4.76 9.09 16.67 38.89 28.57 0.00 11.52 26.15 33.33 57.14 14.29 92.86 7.14
qwen2.5-coder-instruct 7B 0.00 4.26 10.17 0.00 0.00 0.00 5.71 2.99 13.64 11.11 4.76 0.00 0.00 11.90 19.44 28.57 0.00 4.24 30.77 20.00 7.14 0.00 0.00 0.00
qwen2.5-coder-instruct 32B 7.69 6.38 15.25 10.00 28.57 0.00 15.71 4.96 15.15 27.78 33.33 4.76 36.36 10.71 47.22 42.86 2.94 10.91 33.85 20.00 92.86 28.57 71.43 28.57
star-coder-instruct 15B 15.38 4.26 10.17 30.00 42.86 0.00 34.29 3.46 7.58 27.78 0.00 0.00 45.45 16.67 41.67 57.14 0.00 16.97 38.46 26.67 42.86 0.00 42.86 7.14
deepseek-coder-instruct 7B 7.69 2.13 10.17 0.00 0.00 0.00 8.57 2.82 5.30 5.56 4.76 0.00 18.18 9.52 16.67 28.57 0.00 4.24 18.46 33.33 21.43 0.00 21.43 0.00
deepseek-coder-instruct 33B 0.00 6.38 6.78 20.00 7.14 0.00 2.86 3.23 6.82 13.89 25.40 1.59 0.00 11.90 19.44 42.86 0.00 10.30 40.00 26.67 14.29 7.14 14.29 0.00
@@
llama3.1-sft 8B 0.00 6.38 15.25 20.00 0.00 0.00 41.43 34.57 26.52 27.78 90.48 17.46 36.36 9.52 11.11 28.57 2.94 0.00 13.85 0.00 35.71 0.00 57.14 0.00
llama3.1-ultrachat 8B 0.00 8.51 5.08 0.00 0.00 0.00 0.00 0.07 7.58 0.00 1.59 0.00 0.00 4.76 22.22 14.29 0.00 4.85 9.23 6.67 14.29 7.14 0.00 0.00
llama3.1-ultrachat-fm 8B 0.00 10.64 15.25 30.00 7.14 0.00 42.86 39.86 31.06 38.89 96.83 17.46 18.18 9.52 13.89 28.57 11.76 0.00 10.77 0.00 42.86 7.14 64.29 7.14
llama3.1-tulu 8B 0.00 6.38 1.69 0.00 0.00 0.00 0.00 1.43 6.06 0.00 3.17 0.00 0.00 13.10 30.56 28.57 0.00 6.67 21.54 13.33 21.43 7.14 21.43 0.00
llama3.1-tulu-fm 8B 0.00 6.38 11.86 30.00 7.14 0.00 44.29 40.20 32.58 38.89 98.41 17.46 18.18 13.10 11.11 42.86 2.94 4.24 9.23 0.00 71.43 7.14 64.29 0.00
qwen2.5-sft 7B 0.00 4.26 13.56 10.00 7.14 0.00 40.00 32.77 30.30 27.78 88.89 17.46 45.45 10.71 19.44 28.57 2.94 1.21 12.31 0.00 42.86 0.00 35.71 0.00
qwen2.5-coder-sft 7B 0.00 8.51 22.03 30.00 7.14 0.00 44.29 40.81 35.61 36.11 98.41 17.46 54.55 10.71 19.44 42.86 8.82 4.24 20.00 13.33 64.29 0.00 57.14 0.00
deepseek-coder-sft 7B 0.00 4.26 16.95 0.00 0.00 0.00 35.71 30.70 36.36 25.00 85.71 15.87 54.55 13.10 16.67 28.57 2.94 4.24 9.23 6.67 14.29 7.14 28.57 7.14
@@ Appendix F Prompt Design
Prompt for ProofInfill Task (lang: a placeholder to be replaced by each formal specification language name.) # Task Description (ProofInfill) 1. Please fill in the [MASK] in the following formal proof in formal specification language lang according to the given requirement. 2. Please fill in the [MASK] in the following formal proof in lang according to the given requirement. 3. Please complete the given formal proof in lang following the requirement below by filling in the [MASK]. 4. Please fill in the [MASK] in the following formal proof in lang according to the requirements below. 5. Please fill in the [MASK] in the following formal proof in lang according to the given requirement. (Randomly choose one of the above.) You only need to return the completed lang formal specification (together with the provided formal specification) without explanation. # Input
@@
Prompt for Code2Proof Task # Task Description (Code2Proof) 1. Please fill in the [MASK] in ACSL according to the given requirement and ACSL specification. 2. Please fill in the [MASK] in ACSL according to the given requirement. 3. Please fill in the [MASK] in ACSL according to the given ACSL specification. 4. Please fill in the [MASK] in ACSL according to the given requirement and ACSL specification. 5. Please infill the [MASK] in ACSL according to the given requirement. (Randomly choose one of the above.) You only need to return the completed ACSL formal specification (together with the provided formal specifications and C programs) without explanation. # Input